OpenStudy (anonymous):

To the nearest .001, lim (1-x)^(1/x) x->0+ how to solve?

OpenStudy (anonymous):

easy use the natural log. e^ln((1-x)^(1/x)) now use the rules of ln :) = e^ln((1/x)ln(1-x)) = (1/x)ln(1-x) lim (1/x)ln(1-x) x -> +0 are you allowed to use L'H rule?

OpenStudy (anonymous):

what is the l'h rule?

OpenStudy (anonymous):

l'hospital's rule

OpenStudy (anonymous):

it is used to solve problems like these because you have 0 and infinity

OpenStudy (anonymous):

we never used the natural log for this question though

OpenStudy (anonymous):

well the answer to that question is .368 but what im wondering is how did it end up with that answer

OpenStudy (anonymous):

are you going ahead of your class or just learning this for the hell of it?

OpenStudy (anonymous):

this is what i learned in class but i have no idea the process of it

OpenStudy (anonymous):

do you understand the first steps?

OpenStudy (anonymous):

of your method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have to know this for a midterm so I don't mind helping you understand it infact I would be happy to help

OpenStudy (anonymous):

thanks, and i understand that your useing th natural log in order to bring down the exponent

OpenStudy (anonymous):

but how did the answer end up to .368?

OpenStudy (anonymous):

right that is only the first step you need to use L'H rule to solve then you have to sub it back into the original problem

OpenStudy (anonymous):

its a little more complex than what I have posted if you notice I have infinity(0)

OpenStudy (anonymous):

if you sub in 0+

OpenStudy (anonymous):

I haven't gone through the method entirely do you want me to?

OpenStudy (anonymous):

yes please do!

OpenStudy (anonymous):

so we need to apply L'H rule which is used to solve problems such as these (remember to denote that you are using this rule or they will probably doc you) I will do it on paper for you so give me some time

OpenStudy (anonymous):

I will show you how to denote it

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

to set up infinity/infinity or 0/0 we take one of the functions and divide it so f(x)g(x) f(x) = infinity g(x) = 0 so we can go f(x)(1/g(x)) for 0/0 or g(x)/(1/f(x)) for infinity/infinity then we take the derivative and simplify :)

OpenStudy (anonymous):

usually it is best to go infinity/infinity tbh

OpenStudy (anonymous):

if you dont know but it is usually pretty obvious which functions will derive nicely and which wont

OpenStudy (anonymous):

alright, so what is the next step?

OpenStudy (anonymous):

Im working on it

OpenStudy (anonymous):

im doing it the wrong way first to show you the two choices

OpenStudy (anonymous):

if thats ok

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

see that if I keep differentiating this it will always be 0/0 and thus it is the wrong choice

OpenStudy (anonymous):

you can keep applying L'H rule over and over until you end up with a final value but only one choice will work to my knoweldge

OpenStudy (anonymous):

I think I might of made a mistake deriving 1/x but yeah you see how that choice doesn't work I will actually solve it now

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

do you kind of see what I'm doing now though?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

there is one more step let me check my notes

OpenStudy (anonymous):

yup :) I got this pellet so we know that everything is equal to -1 so e^(u) = the solution -1 = u

OpenStudy (anonymous):

e^(-1) = 0.368

OpenStudy (anonymous):

do you follow how I solved that?

OpenStudy (anonymous):

because ln((1-x)^(1/x)) = -1 we are just left with e^(-1)

OpenStudy (anonymous):

ta da

OpenStudy (anonymous):

at the 2nd equal. on top, how did you get the -1?

OpenStudy (anonymous):

2nd equal sign. the -1 on top of the 1-x

OpenStudy (anonymous):

take the derivative of ln(1-x) and then take the derivative of x

OpenStudy (anonymous):

the derivative of ln(1-x) = x'/x so whats the derivative of 1-x

OpenStudy (anonymous):

its -1/1-x

OpenStudy (anonymous):

ohhh okayy

OpenStudy (anonymous):

then the -1

OpenStudy (anonymous):

I hope you understand L'H rule now if you have any other questions ask away

OpenStudy (anonymous):

yeah, is there an easier way to do this problem? i dotn remember being this long

OpenStudy (anonymous):

Simplify always before moving on to the next derivative, unless it is too much of a mess to simplify

OpenStudy (anonymous):

this is the only way to my knowledge and it really isn't that difficult if you have a firm grasp of derivatives and algebra. I recommend you practice some questions. here is a tip look at the two functions and put the cleanest derivative on the bottom

OpenStudy (anonymous):

the reason I took so long is because I was explaining how to do this as well, and also I'm half asleep

OpenStudy (anonymous):

trust me, your not the only one hahahha

OpenStudy (anonymous):

Seriously though this pellet is easy try the question lm x^(x^(2)) x->0+

OpenStudy (anonymous):

or don't it is up to you I have the answer from my prof but if you get stuck at a part I can help you through it

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

is it always the natural log rule?

OpenStudy (anonymous):

that is the only way I know how to solve this type of problem.

OpenStudy (anonymous):

okay i dont know after i got x^2(ln x)

OpenStudy (anonymous):

or did do that wrong

OpenStudy (anonymous):

oh wait, now i find the dirivatives of both right?

OpenStudy (anonymous):

this is a bad example lol you dont need to use L'H rule sorry about that. You can just solve it now lol

OpenStudy (anonymous):

and e^(0) = 1

OpenStudy (anonymous):

oh.....

OpenStudy (anonymous):

I will give you a problem actually involving L'H rule this time I promise lol

OpenStudy (anonymous):

lim (cos(x))^(1/(x^(2))) x-> 0+

OpenStudy (anonymous):

this is the problem that requires L'H rule sorry about that

OpenStudy (anonymous):

oh god this looks complicated

OpenStudy (anonymous):

its not too bad

OpenStudy (anonymous):

just use the same method I did

OpenStudy (anonymous):

okay. 1/(x^2)(ln cos x)?

OpenStudy (anonymous):

now i find the derivative of the 1/x^2 and ln cos x

OpenStudy (anonymous):

right because 1/0+ is infinity and ln(cos(0)) = ln(1) = 0

OpenStudy (anonymous):

so now you have two functions g(x)f(x) here is where you apply the rule first look at the two functions which one will be the neatest to derive?

OpenStudy (anonymous):

cox x?

OpenStudy (anonymous):

no cos(x) is just going to be a repeated derivative and will never completely simplify so we don't want to put it as the denominator

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

1/infinity = 0 so we are going for 0/0

OpenStudy (anonymous):

so put 1/x as the denominator

OpenStudy (anonymous):

I would recommend writing it down exactly how I wrote it down.

OpenStudy (anonymous):

how does it go?

OpenStudy (anonymous):

not you have to apply the rule multiple times in this question and remember that tan(x) = sin(x)/cos(x)

OpenStudy (anonymous):

sorry, its kinda hard doing it on the comp. without writing it down

OpenStudy (anonymous):

write it down lol dont do it on the computer

OpenStudy (anonymous):

you gave up fair enough Have a good night :)

OpenStudy (anonymous):

the answer was 1/(e)^(1/2)